Tyler has two savings accounts that his grandparents opened for him. One of the accounts pay 6% annual interest, whereas the other pays 8% annual interest. If you have $800 more invested at 8% than you invested at 6%, how much do you have invested in each account if the total amount of interest you earn in a year is $120?
Just need to know how to set up the problems

Answer by macston(5194)

.
S = amount at 6%; E = amount at 8% = S+$800
.
0.06S + 0.08E = $120
That is all you need to know to set it up.
.
The following is given in case you have trouble:
0.06S + (0.08)(S+$800)=$120
0.06S+0.08S+$64=$120
0.14S=$56
S=$400
ANSWER 1: You have $400 in the 6% account.
.
E=S+$800=$400+$800=$1200
ANSWER 2: You have $1200 in the 8% account.
.
CHECK:
0.06S+0.08E=$120
0.06($400)+0.08($1200)=$120
$24+$96=$120
$120=$120
.